OpenStudy (anonymous):

Simplify the equation and write it in standard form. If the equation contains decimals, multiply each term by a constant that will make all coefficients integers. Example: 0 = x3 + 2.75x2 – 6.25 If this equation is multiplied by 4, the decimals will become integers. 4(0 = x3 + 2.75x2 – 6.25) 0 = 4x3 + 11x2 – 25 My equation: 168=(x+1)(x-4)(x)

OpenStudy (anonymous):

@GabeBae

OpenStudy (gabebae):

What type of math is this?

OpenStudy (anonymous):

algebra 2. This is what i had, but my teacher said it was wrong: 168=(x+1)(x-4)(x) 168(x^2-4x+x-3)(x) 168(x^2-3x-4)(x) V=x^3-3x^2-4x-168

OpenStudy (anonymous):

She said this 168=(x^2-4x+x-3)(x) is not equal to V=x^3-3x^2-4x-168

OpenStudy (gabebae):

What's the question that's being asked her?

OpenStudy (anonymous):

Simplify the equation and write it in standard form. If the equation contains decimals, multiply each term by a constant that will make all coefficients integers.

OpenStudy (anonymous):

@GabeBae

OpenStudy (gabebae):

meh

OpenStudy (anonymous):

do you know what to do

OpenStudy (gabebae):

Not really

OpenStudy (gabebae):

I only took algebra and i'm taking geometry this year

OpenStudy (anonymous):

you havent taken algebra 2?

OpenStudy (gabebae):

no

OpenStudy (anonymous):

@TayTay23

OpenStudy (anonymous):

Can I come back in a couple hours? I won't have time to answer right now.

OpenStudy (anonymous):

@TayTay23 yes

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

@freckles